Tarzan and Jane adopted a new born monkey and called him Bongo. Unlike other monkeys, Bongo likes dried strawberries and consumes 20 strawberries per week. The nearest strawberry producer to Tarzan and Jane’s house (tree in fact) is far away. Tarzan and Jane estimate that they’ll have to use a day of their neighbor’s elephant every time they shop from that farmer. The neighbor charges them $40 for each trip. One strawberry costs $4. Jane also calculates the holding cost of 1 strawberry as $1 per week, because strawberries occupy a lot of place in their tree. Assume that Tarzan and Jane will never let Bongo be hungry (i.e., no backorders).

P.S. In Tarzan’s forest, each week has 5 days.

When should Tarzan and Jane place an order (i.e., calculate the ROP)?
Answer #1

There are 5 days in a week and the average weekly demand is 20 strawberries

For 5 days the demand is 20 strawberries. So for one day demand = 20/5 = 4 strawberries

Therefore average daily demand (d) = 4 strawberries

Lead time(L) = 1 day

Reorder point = d × L = 4 × 1 = 4 strawberries

So Tarzan and jane should place an order whenever the stock drops to 4 strawberries